Calculate the reaction entropy (S°rxn) for the following reaction: 3 NO2 + H2O 2 HNO3 + NO


1
Expert's answer
2022-05-11T03:09:03-0400

ΔS°rxn = ΔS(NO) + 2ΔS(HNO₃) - 3ΔS(NO₂) - ΔS(H₂O)

ΔS°rxn= 210.65 + (2 × 155.6) - (3 × 239.9) - 69.94 = 210.65 + 311.2 - 719.7 - 69.94 = 521.85 - 789.64 = -267.79 J/K (decrease in entropy)